OpenStudy (anonymous):

Bob invested $23,000, part at 16% and part at 15%. If the total interest at the end of the year is $3500, how much invested at 16%

OpenStudy (matt101):

Let x be the amount invested at 16% Let y be the amount invested at 15% That means: x + y = 23000 0.16x + 0.15y = 3500 Try to solve it from there :)

OpenStudy (anonymous):

Possible answers are 4,000, 5,000, 6,000, 18,000

OpenStudy (anonymous):

x + y = 23000 0.16x + 0.15y = 3500 0.16x + 0.16y = 23000*0.16 0.16x + 0.15y = 3500 ================== Subtracting, you get: 0.01y = 23000*0.16 - 3500 => y = (23000*0.16 -3500)/0.01 => x = 23000 - (23000*0.16-3500)/0.01

OpenStudy (anonymous):

If you calculate those values, you get x = 5000.

OpenStudy (anonymous):

Thank You : D